UT20090613 Telescope Performance Verification
with IRTC at LFBG

Observers: DThompson, MPedani (LBT)
Telescope Operator: DGonzalez-Huerta
Software Support" TEdgin (Tucson)
Telescope Support: DOfficer

Summary

  • Lost 2h to MCR problem. Bypassed but not fixed.
  • Guider gain factor tuning - recommend ~0.75 for now
  • Attempted collimation model obs, interrupted by clouds
  • Lost last 2.5h to heavy clouds.

Details (Time in UTC)

01:27:30.613 Closed-dome ACQUIRE mode preset...OSS hung on collimating M3. Stop and restart OSS.
01:32:01.012 Closed-dome ACQUIRE mode preset...successful.
01:33 Toutside=10.4 Tmirror=4.9 RH=29.7% Wind ~14m/s with gusts to 18.5.

02:40 Open. Some thin cirrus at sunset. Winds around 15m/s. MCR problem with the telescope.

04:40 Problem with MCR identified and bypassed. Ready to observe.

04:40:15.873 pointto WT10_228 TRACK mode . IE -99 and CA 13. Way out of focus.

04:47:00.653 TRACK...ooops.
04:48:02.069 ACTIVE to collimate on BS9160. WFS: 3398, 2509, 2509, 1803, 768, 383, 253, 438, 242, 210

irc OffsetPointing 0.0 12.33 7.785 RADEC MOUNT true ABS left

takeseq 2.0 30 for sky...045929 0001_cube
offset 8.0 9.0
takeseq 2.0 30 for sky...050135 0002_cube

05:10 GUIDE FACTOR TEST

N1 = First Guider Image N2 = Last Guider Image Single = First IRTC Image irtc.20090613.* of the sequence

Note: we take Sequences of 100 IRTC images 2s each.

GFact N1 N2 Single Cube Notes
0.75 0424 0514 051122 3  
1.00 0525 0597 051604 4  
0.75 0608 0679 052014 5  
1.00 0688 0765 052411 6  
0.75 0771 0848 052809 7  
1.00 0857 0927 053224 8  
0.75 0934 1004 053637 9  
1.00 1013 1082 054037 10  
0.75 1093 1158 054448 11  
1.00 1181 1254 054940 12  
0.75 1699 1769 061609 13  
1.00 1775 1855 062014 14  
0.75 1867 1941 062515 15  
1.00 1952 2019 063000 16  

06:46 Check Pointing WT10_226

Now CA= 9.0 IE = -115

06:50 Rotator hangs

06:53 IIF Preset command timed out in 300 seconds for left IRTC mode TRACK from IIF

06:58 Collimation DATA

irtc.20090613.070202.fits

Test image, remove split and go again.

NOW Start TAKING COLLIMATION DATA

Star # Image name EL
1 irtc.20090613.070641.fits 71
2 irtc.20090613.071416.fits 77
3 irtc.20090613.071937.fits 82
4 irtc.20090613.072937.fits 85
5 irtc.20090613.073456.fits 76
6 irtc.20090613.073917.fits 70
7 irtc.20090613.074344.fits 62
8 irtc.20090613.075219.fits 55
9 irtc.20090613.083007.fits 52
10 irtc.20090613.083835.fits 37

9:00 Clouds getting heavier. Closing dome for now. Toutside=6.7, Tmirror=7.0, RH=53% and rising, Wind 8 with gusts to 15 m/s.

10:06 Call it a night.

-- MarcoPedani - 13 Jun 2009
I Attachment Action Size Date Who Comment
daq.20090613.gifgif daq.20090613.gif manage 39 K 08 Jul 2009 - 18:28 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r8 - 08 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback